OpenStudy (dmezzullo):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (anonymous):

2x-5=100

OpenStudy (dmezzullo):

show the steps plz

OpenStudy (anonymous):

2x-5=100 2x=95 x=95/2

OpenStudy (dmezzullo):

u sure

OpenStudy (dmezzullo):

omg ive done skewl all day my brain is sooo freakin fried and i wanna tlk to my gf

OpenStudy (anonymous):

dmezzullo, you really need to try and understand this. If we just give you the asnwers it won't do you any good on a test.

OpenStudy (dmezzullo):

i try to understand it

OpenStudy (anonymous):

If we have an equation x = 2 What is x?

OpenStudy (dmezzullo):

value not known

OpenStudy (anonymous):

Yes it is! Since the EQUAL sign means they are EQUAL! So x is 2